OpenStudy (anonymous):

Ivan estimated the volume of his bathtub to be 31gal . The actual volume of his bathtub was 37gal. Find the absolute error and the percent error of Ivan's estimate. If necessary, round your answers to the nearest tenth

OpenStudy (anonymous):

is percent error 60.3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@slimjim77 When you find absolute error, it is: (absolute value of the difference between actual and esimated volume) divided by actual volume = (37-31)/37 = 6/37 = 16.2%

OpenStudy (anonymous):

Absolute Error is equal to the absolute value of the difference of the estimated and actual. In this case, it is 6 gal. Percent Error is equal to the Absolute Error divided by the Actual value. In this case, it is 16.21%.

OpenStudy (anonymous):

ninja'd by easyaspi314.

OpenStudy (anonymous):

question asked for nearest tenth..so I had 16.2%

OpenStudy (anonymous):

Thanks both of you

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

I think easyaspi314 deserves the medal because he answered the question first.

OpenStudy (anonymous):

I dont need medal. Give medal to @ColossalMagnitude

OpenStudy (anonymous):

I'm not here for medals.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

well both of you got medals lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Well, that's certainly nice of you easyaspi, thanks, lol.

OpenStudy (anonymous):

hey i have one more question

OpenStudy (anonymous):

Kira is estimating the height of her office building. The actual height of her office building is 10.4 . Kira's estimate is 11 . Find the absolute error and the percent error of Kira's estimate. If necessary, round your answers to the nearest tenth.

OpenStudy (anonymous):

so i did (11-10.4)/11-0.6/11

OpenStudy (anonymous):

its not correct and im not sure wats wrong

OpenStudy (anonymous):

Why did you do 11 - 0.6? You should have done .6 / 10.4.

OpenStudy (anonymous):

Which is about 5.8%.

OpenStudy (anonymous):

You have to divide by the ACTUAL measurement. So it is .6/10.4 = 5.8%

OpenStudy (anonymous):

alright i see whats wrong now thanks

OpenStudy (anonymous):

You sure you understand?

OpenStudy (anonymous):

was the rest of it correct

OpenStudy (anonymous):

@ColossalMagnitude gets medal, for sure! lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

(12-11)/12=1/12

OpenStudy (anonymous):

Diane estimated the length of a room is her house to be 12ft . The actual length of the room is 11ft .

OpenStudy (anonymous):

so it would be 8.3 percent?